OpenStudy (anonymous):

The slope of a line is 1, and the y-intercept is -1. What is the equation of the line written in slope-intercept form? y = x - 1 y = 1 - x y = -x - 1 C i got C

OpenStudy (anonymous):

@texaschic101

OpenStudy (texaschic101):

y = x - 1 look at it this way... y = mx + b m is your slope and b is your y inyercept

OpenStudy (anonymous):

what i did was y=-1x-1......then i simplified to y=-x-1

OpenStudy (anonymous):

y=mx+b y=(1)*(x)+(-1)

OpenStudy (texaschic101):

your slope is not -1, it is positive 1

OpenStudy (anonymous):

oh oops i wrote down wrong lol youre write thanks again

OpenStudy (texaschic101):

anytime :)